# Large Language Models and Code Security: A Systematic Literature Review

### v2

@@ V-B Issues in Using LLMs for Vulnerability Detection
While LLMs show considerable promise for vulnerability detection, several challenges and limitations persist. In this section, we aim to unwrap the limitations highlighted by the literature. Table IV summarizes the challenges faced by LLMs in detecting vulnerabilities in code, as identified by the reviewed studies. In this table, we indicate with a green checkmark () which challenge was highlighted in which study. Notably, most studies focused on the prevalence of false positives, discussed in 7 papers, and challenges with complex tasks, highlighted in 6. In contrast, the other three challenges were mentioned in only one or two studies out of 12.
@@
TABLE IV: Challenges in using LLMs for vulnerability detection identified in the studies
@@
References False Positives Context Awareness Complex Tasks Verbosity and Non-Compliance Token Limits
Steenhoek et al. [52]
Happe et al. [53]
Yu et al. [54]
Ullah et al. [55]
Purba et al. [56]
Liu et al. [57]
Ding et al. [58]
Çetin et al. [59]
Pelofske et al. [60]
Chen et al. [61]
Sun et al. [62]
Gao et al. [63]
Zhang et al. [46]
Li et al. [64]
Hu et al. [65]
Zhou et al. [66]
Tamberg et al. [67]
Total number of studies 11 3 6 1 2
False positives are one of the key challenges for LLMs, where the model incorrectly flags a non-existent vulnerability.
@@
Various works highlighted the challenge faced by LLMs with false positives, when dealing with vulnerability detection. Ding et al. [58] reported that commonly used benchmarks often overestimate model performance, with a state-of-the-art 7B model achieving a 68.26% F1 score on a popular dataset BigVul, but only a 3.09% on the dataset they proposed. Similarly, Çetin et al. [59] found that GPT-4, despite being the top-performing model in their study, still had a false positive rate of 63%, with other models performing even worse, reaching up to 97%. Pelofske et al. [60] revealed that, while some models achieved high recall and precision rates (up to 100%) for certain vulnerability types, they still exhibited significant rates of false positives. Chen et al. [61] also noted that while ChatGPT was able to identify vulnerabilities in blockchain smart contracts, it struggled with precision, leading to false positives. Zhou et al. [66], Tamberg et al. [67], and Purba et al. [56] found that, although LLMs outperform traditional vulnerability detection approaches in identifying vulnerabilities, they also exhibit high false positive rates. We further discuss this aspect in Subsection V-C.
@@
Some works have looked into proposing solutions to address this challenge. Li et al. [64] noticed that LLMs often produce false positives when used as static code analyzers. To address this, they proposed a method that prompts ChatGPT to summarize functions, breaking down tasks step by step and progressively refining results to improve accuracy. Their approach successfully eliminated 16 out of 20 false positives. Hu et al. [65] addressed the false positives issue by proposing a framework called GPTLens, which splits the conventional one-stage detection into two phases: generation and discrimination. In this framework, an LLM plays a dual role. In the generation phase, it acts as the Auditor, aiming to generate a wide range of vulnerabilities to capture the correct ones. In the discrimination phase, it acts as the Critic, evaluating these vulnerabilities to reduce false positives. Empirical results show that GPTLens significantly improves vulnerability detection over traditional methods.
@@
Context awareness is another challenge that LLMs face when detecting vulnerabilities in code. Happe et al. [53] pointed out that hallucinations stem from a lack of contextual understanding and grounding. Sun et al. [62] addressed this issue by developing GPTScan, which enhances context awareness by combining GPT with static analysis techniques. GPTScan parses code, identifies potential vulnerabilities, and maps them to known types, significantly improving detection precision and recall.
@@
Some researchers have reported that LLMs often struggle with more complex tasks, which often include real-world scenarios. For instance, Yu et al. [54] found that while GPT-4 outperformed other models in security code review, its performance was still limited in more complex scenarios. Ullah et al. [55] also highlighted that LLMs, including advanced models like GPT-4, exhibited unpredictable responses and incorrect reasoning when tasked with identifying security vulnerabilities in real-world scenarios. Moreover, Purba et al. [56] noted that LLMs performed well in detecting certain vulnerability patterns, but still could not surpass traditional tools in handling complex tasks, such as SQL injection and buffer overflow detection. Liu et al. [57] developed a benchmark to assess LLMs on progressively difficult tasks. They revealed that LLMs perform well (over 80% accuracy) at simpler tasks, such as identifying and classifying vulnerabilities. However, their accuracy dropped (below 30%) when faced with more complex tasks, such as detecting the root causes of vulnerabilities across extensive codebases. Finally, Ding et al. [58] observed that LLMs are not able to meet the requirements of effective vulnerability detection in real-world scenarios.
@@
Yu et al. [54] found that, while GPT-4 was effective in detecting security vulnerabilities and outperformed other models, it faced verbosity and non-compliance challenges. These issues emerge when the model occasionally provides overly detailed responses or fails to meet prompt requirements, respectively. Both issues can affect the quality and trustworthiness of responses, in general.
@@
Chen et al. [61] discussed the challenge of token limits, when ChatGPT is used for identifying vulnerabilities in blockchain smart contracts. They found that the maximum code sizes that GPT-4 and GPT-3.5 models could effectively analyze are around 33 KB and 16 KB, respectively. To address similar challenges posed by input length limitations, Zhang et al. [46] proposed a solution based on sliding windows that divide long sequences into overlapping sequences of tokens. In addition, they introduced a right-forward embedding strategy, which enhances context by making each token aware of tokens that come after.
@@
Despite the challenges we have highlighted so far, most of the studies we have covered agree that LLMs have the potential to become effective tools for static analysis, provided additional fine-tuning and context-aware enhancements.
V-C Comparison of LLMs with Traditional Vulnerability Detection Approaches
@@
Table V, and Table VI provide an overview of the different LLM models and traditional analysis tools that have been used in each study, for comparing performance. Table V lists the papers where LLMs showed better vulnerability detection capabilities than traditional approaches. In contrast, Table VI highlights the papers where traditional approaches performed better, identifying a higher percentage of vulnerabilities. From the sheer number of papers included in each table, it is clear that, in most literature works, LLMs showed greater potentiality in discovering security vulnerabilities.
@@
TABLE V: Studies showing that LLMs outperformed traditional approaches, based on the number of vulnerabilities detected.
@@
Yu et al. [54] GPT (3.5, 4, 4 Turbo), Llama 2 (7B, 70B), Gemini Pro Bandit, CodeQL, CppCheck, Semgrep, SonarQube
Zhang et al. [68] ChatGPT-4 Bugram, CFGNN
Mohajer et al. [44] ChatGPT (3.5 Turbo, 4) Infer
Thapa et al. [45] GPT (2, J, MegatronGPT-2), BERT, MegatronBERT, DistilBERT, RoBERTa, CodeBERT BiLSTM, BiGRU
Noever et al. [69] GPT-4 Snyk, Fortify
Wen et al. [47] Code-llama-7b, GPT-3.5 Devign, Reveal
Omar et al. [48] GPT-2 SySeVR
Gao et al. [63] GPT (3.5, 4), Llama 2 (7b-chat-hf, 13b-chat-hf, 70b-chat-hf), Vicuna (7b-v1.5, 7b-v1.5-16k, 13b-v1.5, 13b-v1.5-16k, 33b-v1.3), ChatGLM2-6b, Baichuan2-13B-Chat, Internlm-20b-chat, CodeLlama-34b-Instruct, Falcon-40b-instruct, Platypus2-70B-instruct LineVul, CodeXGLUE, Cppcheck, Flawfinder, BinAbsInspector
Wen et al. [51] CodeBERT, CodeT5, UniXcoder, EPVD, LineVul, SVulD Cppcheck, Flawfinder, RATS, Semgrep, Devign, Reveal
Hu et al. [65] GPT-4 Slither
Ozturk et al. [70] ChatGPT-3.5 Betterscan CE, OWASP WAP, phpcs-security-audit (PHPcs), Pixy, Progpilot, Psalm, RATS, RIPS, SonarQube, Visual Code Grepper (VCG), Horusec
Purba et al. [56]* GPT-3.5-Turbo, Davinci, CodeGen-2B-multi model Checkmarx, Flawfinder, RATS
Zhou et al. [66]* CodeBERT, G-BERT, CodeT5, UniXcoder, StarCoder, CodeLlama, Mistral, DSCoder, StarCoder2, CodeQwen, LLama3, Phi3 Bandit, CodeChecker, CodeQL, Contrast Codesec Scan, Cppcheck, DevSkim, Dlint, Flawfinder, Graudit, Horusec, Infer, Insider, Semgrep, SonarQube, SpotBugs
Tamberg et al. [67]* GPT (4, 4-turbo), Claude 3 Opus CodeQL, SpotBugs
@@
TABLE VI: Studies showing that traditional approaches outperformed LLMs, based on the number of vulnerabilities detected.
@@
Bakhshandeh et al. [71] GPT-3.5-Turbo Bandit, Semgrep, SonarQube
@@
Zhang et al. [46] GPT (3.5, 4), Llama 2, CodeLlama, WizardCoder, CodeBERT, GraphCodeBERT, PLBART, CodeT5, CodeGen VulTeller, Transformer
@@
Chen et al. [61] GPT (3.5, 4) Conkas, Oyente, Smartcheck, Honeybadger, Maian, Securify, Slither, Osiris, Smartian, Confuzzius, Sailfish, Mythril, Solhint, AChecker
Ni et al. [72] ChatGPT-3.5 Turbo Devign, ReVeal, IVDetect, LineVul, SVulD
According to Yu et al. [54], traditional code analysis tools, despite their long-standing use, often struggle with issues such as high false positive rates, limited input ranges, and poor scalability. In contrast, LLMs, particularly GPT-based models, provide higher detection rates, compared to contemporary SATs. While most studies suggest that LLMs outperform traditional approaches in vulnerability detection, LLMs often exhibit also high false positive rates. For instance, Purba et al.[56] found that LLMs achieved higher detection rates than SATs, but exhibited false positive rates exceeding 60%, compared to rates under 45% for SATs like Flawfinder [73] and Checkmarx [74].
Building on this, Zhou et al. [66] found that LLMs can detect up to 100% of code vulnerabilities, but they often produce false positive results. In contrast, although SATs achieve lower vulnerability detection rates of up to 44.4%, they also exhibit lower false positive rates. The authors concluded that their extensive comparison of 15 different SATs, against 12 state-of-the-art LLMs, revealed significant differences in detecting software vulnerabilities across repositories of three programming languages: Java, C, and Python. Similarly, Tamberg et al. [67] highlighted that LLMs can identify more vulnerabilities than traditional SATs, but they are prone to non-deterministic behavior and higher false positive rates. Despite these challenges, LLMs adaptability and contextual understanding make them effective tools for vulnerability detection, especially in identifying a wider range of security issues.
Ozturk et al. [70] designed a benchmark of 92 web application vulnerabilities, with a focus on PHP. Compared to 11 different SATs, ChatGPT was able to identify approximately 62-68% of the security vulnerabilities, while the best-performing SAT detected only 32%. However, ChatGPT exhibited a very high false positive rate of 91%, while the highest false positive rate for the SATs was 82%.
Bakhshandeh et al. [71] investigated the applicability of ChatGPT (in the form of GPT-3.5), for vulnerability detection on Python code. They compared the performance of ChatGPT against three SATs In the comparison, ChatGPT was used not only as a stand-alone tool but also as an assistant module for the SATs. The study indicated that while LLMs do not outperform traditional SATs, they show potential when combined with SATs to reduce false positives and negatives.
Further research by Chen et al. [61] explored the performance of ChatGPT against 14 smart contract vulnerability detection tools. Their results showed that ChatGPT performed worse than these tools on 71.4% (5/7) of vulnerabilities, showing high efficiency but more detection failures than 71.4% (10/14) of the tools tested. However, in a related experiment by Hu et al. [65], the LLM framework outperformed the smart contract detection tool Slither [75].
Other researchers turned to ML-based approaches as baselines to assess the effectiveness of LLMs in detecting code vulnerabilities. For instance, Gao et al. [63] employed three Deep Learning (DL) models, along with three rule-based SATs as baselines in their research. Their results suggest that GPT models outperform SATs and other DL models, such as VulBERTa and CodeXGLUE in vulnerability detection. However, open-access LLM models, (e.g., Llama 2 and Vicuna), do not achieve the same performance, even when trained on the same datasets. Overall, there seems to be a significant gap in performance, between advanced models such as GPT-3.5 and large-scale open-access alternatives. Similarly, Zhang et al. [68] concluded that ChatGPT has demonstrated significant advantages over ML-based approaches, such as Bugram [76] and CFGNN [77], in vulnerability detection tasks.
These findings align with the conclusions of other researchers, such as Thapa et al. [45], who reported that transformer-based language models outperform recurrent neural network (RNN)-based models. Specifically, transformer-based models show better performance than bidirectional long short-term memory (BiLSTM) models [78] and bidirectional gated recurrent unit (BiGRU) models [79]. These results were observed on datasets featuring buffer and resource management errors, related to API function calls.
Furthermore, the framework built upon ChatGPT by Wen et al. [47], was able to capture more vulnerabilities than the ML tools Devign [80] and Reveal [81]. Similarly, SySeVR, an ML-based model, failed to match the performance of LLMs in vulnerability detection, confirming LLMs superiority [48].
While in the literature, LLMs generally outperform traditional ML approaches, there are notable exceptions. For instance, in a study by Zhang et al. [46], the ML-based model VulTeller [82] and the transformer-based model from VELVET [83] demonstrated superior performance over LLMs in key metrics for vulnerability detection. The superiority was especially evident in zero-shot scenarios, where these models outperformed LLMs.
@@
In another study, Ni et al. [72] compared sequence-based and graph-based models for vulnerability detection. They found that sequence-based models, such as LineVul [84] and SVulD [85], generally outperformed graph-based models like Devign, ReVeal, and IVDetect [86] in accuracy and stability. However, graph-based models showed strengths in certain types of vulnerabilities like API Abuse and Security Features. The study also noted that LLMs are not as effective for direct use in vulnerability detection, and that their performance is highly dependent on prompt design.
@@
Some studies indicate that combining LLMs with SATs can improve vulnerability detection rates. For instance, Li et al. [87] observed that LLMs often struggle with complex code reasoning for vulnerability detection and proposed IRIS, a method that combines LLMs with SATs. IRIS uses LLMs to find specific taint specifications related to a vulnerability class (CWE) in third-party library APIs within a project. IRIS improves these specifications with CodeQL to find security vulnerabilities. They evaluated IRIS on a dataset they created with 120 security vulnerabilities found in real-world Java projects, and validated manually. Results showed that IRIS achieved a 35% increase in vulnerability detection compared to using CodeQL alone, while also reducing the number of false positives by 80%. In summary, the combination of LLMs and SATs has enhanced the analysis of vulnerabilities.
@@
Similarly, Li et al. [88] proposed LLift, a framework that combines LLMs with static analysis (using UBITect, a UBI discovery tool), to improve bug detection in large and complex codebases, with a focus on Use Before Initialization (UBI) bugs in the Linux kernel. First, the pipeline uses UBITect to perform a path-insensitive static analysis for scalability, followed by symbolic execution for precision. Then, in the second phase of the pipeline, an LLM is used to address the undecided cases that UBITect cannot handle. The results showed that LLift reduces the complexity of path-sensitive analysis, improving bug detection accuracy without a high rate of false positives. Additionally, LLift was able to identify 4 UBI bugs in the Linux kernel that UBITect could not confirm.
V-D Factors Influencing LLMs Performance

### v3

@@ 5.2. Issues in Using LLMs for Vulnerability Detection
While LLMs show considerable promise for vulnerability detection, several challenges and limitations persist. In this section, we aim to unwrap the limitations highlighted by the literature. Table 4 summarizes the challenges faced by LLMs in detecting vulnerabilities in code, as identified by the reviewed studies. In this table, we indicate with a green checkmark () which challenge was highlighted in which study. Notably, most studies focused on the prevalence of false positives, discussed in 11 papers. In contrast, the other three challenges were mentioned in only one or two studies out of 17.
@@
Table 4. Challenges in using LLMs for vulnerability detection identified in the studies
@@
References False Positives Context Awareness Complex Scenarios Verbosity and Non-Compliance Token Limits
Steenhoek et al. (Steenhoek et al., 2024)
Happe et al. (Happe and Cito, 2023)
Yu et al. (Yu et al., 2024)
Ullah et al. (Ullah et al., 2024)
Purba et al. (Purba et al., 2023)
Liu et al. (Liu et al., 2024a)
Ding et al. (Ding et al., 2025)
Çetin et al. (Çetin et al., 2024)
Pelofske et al. (Pelofske et al., 2024)
Chen et al. (Chen et al., 2024a)
Sun et al. (Sun et al., 2024)
Gao et al. (Gao et al., 2023)
Zhang et al. (Zhang et al., 2024d)
Li et al. (Li et al., 2023)
Hu et al. (Hu et al., 2023)
Zhou et al. (Zhou et al., 2024c)
Tamberg et al. (Tamberg and Bahsi, 2024)
Total number of studies 11 2 6 1 2
@@
False positives are one of the key challenges for LLMs, where the model incorrectly flags a non-existent vulnerability. Ding et al. (Ding et al., 2025) reported that commonly used benchmarks often overestimate model performance, with a state-of-the-art 7B model achieving a 68.26% F1 score on a popular dataset BigVul, but only a 3.09% on the dataset they proposed. Çetin et al. (Çetin et al., 2024) found that GPT-4, despite being the top-performing model in their study, still had a false positive rate of 63%, with other models performing even worse, reaching up to 97%. Pelofske et al. (Pelofske et al., 2024) revealed that, while some models achieved high recall and precision rates (up to 100%) for certain vulnerability types, they still exhibited significant rates of false positives. Chen et al. (Chen et al., 2024a) also noted that while ChatGPT was able to identify vulnerabilities in blockchain smart contracts, it struggled with precision, leading to false positives. Zhou et al. (Zhou et al., 2024c), Tamberg et al. (Tamberg and Bahsi, 2024), and Purba et al. (Purba et al., 2023) found that, although LLMs outperform traditional vulnerability detection approaches in identifying vulnerabilities, they also exhibit high false positive rates. We further discuss this aspect in Subsection 5.3.
@@
Some works have looked into proposing solutions to address this challenge. Li et al. (Li et al., 2023) noticed that LLMs often produce false positives when used as static code analyzers. To address this, they proposed a method that prompts ChatGPT to summarize functions, breaking down tasks step by step and progressively refining results to improve accuracy. Their approach eliminated 16 out of 20 false positives. Hu et al. (Hu et al., 2023) addressed the false positives issue by proposing a framework called GPTLens, which splits the conventional one-stage detection into two phases: generation and discrimination. In this framework, an LLM plays a dual role. In the generation phase, it acts as the Auditor, aiming to generate a wide range of vulnerabilities to capture the correct ones. In the discrimination phase, it acts as the Critic, evaluating these vulnerabilities to reduce false positives. The results show that GPTLens significantly improves vulnerability detection.
@@
Context awareness is another challenge that LLMs face when detecting vulnerabilities in code. Happe et al. (Happe and Cito, 2023) pointed out that hallucinations stem from a lack of contextual understanding and grounding. Sun et al. (Sun et al., 2024) addressed this issue by developing GPTScan, which enhances context awareness by combining GPT with static analysis techniques. GPTScan parses code, identifies potential vulnerabilities, and maps them to known types, significantly improving detection precision and recall.
@@
Some researchers found that LLMs often struggle with complex scenarios,especially those involving real-world code. For instance, Yu et al. (Yu et al., 2024) reported that although GPT-4 outperformed other models in security code review, its performance remained limited in complex scenarios. Ullah et al. (Ullah et al., 2024) also highlighted that LLMs, including advanced models like GPT-4, exhibited unpredictable responses and incorrect reasoning when identifying security vulnerabilities in real-world scenarios. Purba et al.(Purba et al., 2023) found that while LLMs could detect certain vulnerability patterns, they still could not surpass traditional tools in handling tasks such as SQL injection and buffer overflow detection. Liu et al. (Liu et al., 2024a) developed a benchmark to assess LLMs on progressively difficult tasks. They revealed that LLMs perform well (over 80% accuracy) at simpler tasks, such as identifying and classifying vulnerabilities. However, their accuracy dropped (below 30%) when faced with more complex tasks, such as detecting the root causes of vulnerabilities across extensive codebases. Finally, Ding et al. (Ding et al., 2025) observed that LLMs are not able to meet the requirements of effective vulnerability detection in real-world scenarios.
@@
Yu et al. (Yu et al., 2024) found that, while GPT-4 was effective in detecting security vulnerabilities and outperformed other models, it faced verbosity and non-compliance challenges. These issues emerge when the model occasionally provides overly detailed responses or fails to meet prompt requirements, respectively. Both issues can affect the quality and trustworthiness of responses, in general.
@@
Chen et al. (Chen et al., 2024a) discussed the challenge of token limits when using ChatGPT for identifying vulnerabilities in blockchain smart contracts, noting that GPT-4 and GPT-3.5 can effectively analyze up to  33 KB and 16 KB of code, respectively. To address such input length constraints, Zhang et al. (Zhang et al., 2024d) proposed a solution based on sliding windows that divide long sequences into overlapping sequences of tokens. In addition, they introduced a right-forward embedding strategy, which enhances context by making each token aware of subsequent tokens.
@@
Despite the challenges we have highlighted, most of the studies we have covered agree that LLMs have the potential to be effective tools for static analysis, provided additional fine-tuning and context-aware enhancements.
5.3. Comparison of LLMs with Traditional Vulnerability Detection Approaches
@@
Table 5, and Table 6 provide an overview of the different LLM models and traditional analysis tools that have been used in each study, for comparing performance. Table 5 lists the papers where LLMs showed better vulnerability detection capabilities than traditional approaches. In contrast, Table 6 highlights the papers where traditional approaches performed better, identifying a higher percentage of vulnerabilities. From the sheer number of papers included in each table, it is clear that, in most literature works, LLMs showed greater potentiality in discovering security vulnerabilities.
According to Yu et al. (Yu et al., 2024), traditional code analysis tools, despite their long-standing use, often struggle with issues such as high false positive rates, limited input ranges, and poor scalability. In contrast, LLMs, particularly GPT-based models, provide higher detection rates, compared to contemporary SATs. While most studies suggest that LLMs outperform traditional approaches in vulnerability detection, LLMs often exhibit also high false positive rates. For instance, Purba et al.(Purba et al., 2023) found that LLMs achieved higher detection rates than SATs, but exhibited false positive rates exceeding 60%, compared to rates under 45% for SATs like Flawfinder (fla, [n. d.]) and Checkmarx (che, [n. d.]).
Building on this, Zhou et al. (Zhou et al., 2024c) found that LLMs can detect up to 100% of code vulnerabilities, but they often produce false positive results. In contrast, although SATs achieve lower vulnerability detection rates of up to 44.4%, they also exhibit lower false positive rates. Similarly, Tamberg et al. (Tamberg and Bahsi, 2024) highlighted that LLMs can identify more vulnerabilities than traditional SATs, but they are prone to non-deterministic behavior and higher false positive rates. Despite these challenges, LLMs adaptability and contextual understanding make them effective tools for vulnerability detection, especially in identifying a wider range of security issues.
Ozturk et al. (Ozturk et al., 2023) designed a benchmark of 92 web application vulnerabilities, with a focus on PHP. Compared to 11 different SATs, ChatGPT was able to identify approximately 62-68% of the security vulnerabilities, while the best-performing SAT detected only 32%. However, ChatGPT exhibited a very high false positive rate of 91%, while the highest false positive rate for the SATs was 82%.
Bakhshandeh et al. (Bakhshandeh et al., 2023) investigated ChatGPT effectiveness in detecting vulnerabilities in Python code. ChatGPT was used both as a standalone tool and as an assistant to the SATs. While it did not outperform traditional SATs, the study found it can help reduce false positives and negatives when used in combination.
Chen et al. (Chen et al., 2024a) explored the performance of ChatGPT against 14 smart contract vulnerability detection tools. Their results showed that ChatGPT performed worse than these tools on 71.4% (5/7) of vulnerabilities, showing high efficiency but more detection failures than 71.4% (10/14) of the tools tested. However, in a related experiment by Hu et al. (Hu et al., 2023), the LLM framework outperformed the smart contract detection tool Slither (Feist et al., 2019).
Other researchers turned to ML-based approaches as baselines to assess the effectiveness of LLMs in detecting code vulnerabilities. For instance, Gao et al. (Gao et al., 2023) employed three Deep Learning (DL) models, along with three rule-based SATs as baselines in their research. Their results suggest that GPT models outperform SATs and other DL models, such as VulBERTa and CodeXGLUE in vulnerability detection. However, open-access LLM models, (e.g., Llama 2 and Vicuna), do not achieve the same performance, even when trained on the same datasets. Overall, there seems to be a significant gap in performance, between advanced models such as GPT-3.5 and large-scale open-access alternatives. Similarly, Zhang et al. (Zhang et al., 2024b) concluded that ChatGPT has demonstrated significant advantages over ML-based approaches, such as Bugram (Wen et al., 2015) and CFGNN (Zhang et al., 2023b), in vulnerability detection tasks.
These findings align with the conclusions of other researchers, such as Thapa et al. (Thapa et al., 2022), who reported that transformer-based language models outperform RNN-based models. Specifically, transformer-based models show better performance than bidirectional long short-term memory (BiLSTM) models (Li et al., 2018) and bidirectional gated recurrent unit (BiGRU) models (Cho et al., 2014). These results were observed on datasets featuring buffer and resource management errors, related to API function calls.
Furthermore, the framework built upon ChatGPT by Wen et al. (Wen et al., 2024a), was able to capture more vulnerabilities than the ML tools Devign (Zhou et al., 2019) and Reveal (Chakraborty et al., 2021). Similarly, SySeVR, an ML-based model, failed to match the performance of LLMs in vulnerability detection, confirming LLMs superiority (Omar and Burrell, 2023).
While in the literature, LLMs generally outperform traditional ML approaches, there are notable exceptions. For instance, in a study by Zhang et al. (Zhang et al., 2024d), the ML-based model VulTeller (Zhang et al., 2023a) and the transformer-based model from VELVET (Ding et al., 2022) demonstrated superior performance over LLMs in key metrics for vulnerability detection. The superiority was especially evident in zero-shot scenarios, where these models outperformed LLMs.
@@
Table 5. Studies showing that LLMs outperformed traditional approaches, based on the number of vulnerabilities detected.
@@
Yu et al. (Yu et al., 2024) GPT (3.5, 4, 4 Turbo), Llama 2 (7B, 70B), Gemini Pro Bandit, CodeQL, CppCheck, Semgrep, SonarQube
Zhang et al. (Zhang et al., 2024b) ChatGPT-4 Bugram, CFGNN
Mohajer et al. (Mohajer et al., 2024) ChatGPT (3.5 Turbo, 4) Infer
Thapa et al. (Thapa et al., 2022) GPT (2, J, MegatronGPT-2), BERT, MegatronBERT, DistilBERT, RoBERTa, CodeBERT BiLSTM, BiGRU
Noever et al. (Noever, 2023) GPT-4 Snyk, Fortify
Wen et al. (Wen et al., 2024a) Code-llama-7b, GPT-3.5 Devign, Reveal
Omar et al. (Omar and Burrell, 2023) GPT-2 SySeVR
Gao et al. (Gao et al., 2023) GPT (3.5, 4), Llama 2 (7b-chat-hf, 13b-chat-hf, 70b-chat-hf), Vicuna (7b-v1.5, 7b-v1.5-16k, 13b-v1.5, 13b-v1.5-16k, 33b-v1.3), ChatGLM2-6b, Baichuan2-13B-Chat, Internlm-20b-chat, CodeLlama-34b-Instruct, Falcon-40b-instruct, Platypus2-70B-instruct LineVul, CodeXGLUE, Cppcheck, Flawfinder, BinAbsInspector
Wen et al. (Wen et al., 2024b) CodeBERT, CodeT5, UniXcoder, EPVD, LineVul, SVulD Cppcheck, Flawfinder, RATS, Semgrep, Devign, Reveal
Hu et al. (Hu et al., 2023) GPT-4 Slither
Ozturk et al. (Ozturk et al., 2023) ChatGPT-3.5 Betterscan CE, OWASP WAP, phpcs-security-audit (PHPcs), Pixy, Progpilot, Psalm, RATS, RIPS, SonarQube, Visual Code Grepper (VCG), Horusec
Purba et al. (Purba et al., 2023)* GPT-3.5-Turbo, Davinci, CodeGen-2B-multi model Checkmarx, Flawfinder, RATS
Zhou et al. (Zhou et al., 2024c)* CodeBERT, G-BERT, CodeT5, UniXcoder, StarCoder, CodeLlama, Mistral, DSCoder, StarCoder2, CodeQwen, LLama3, Phi3 Bandit, CodeChecker, CodeQL, Contrast Codesec Scan, Cppcheck, DevSkim, Dlint, Flawfinder, Graudit, Horusec, Infer, Insider, Semgrep, SonarQube, SpotBugs
Tamberg et al. (Tamberg and Bahsi, 2024)* GPT (4, 4-turbo), Claude 3 Opus CodeQL, SpotBugs
@@
Table 6. Studies showing that traditional approaches outperformed LLMs, based on the number of vulnerabilities detected.
@@
Bakhshandeh et al. (Bakhshandeh et al., 2023) GPT-3.5-Turbo Bandit, Semgrep, SonarQube
@@
Zhang et al. (Zhang et al., 2024d) GPT (3.5, 4), Llama 2, CodeLlama, WizardCoder, CodeBERT, GraphCodeBERT, PLBART, CodeT5, CodeGen VulTeller, Transformer
@@
Chen et al. (Chen et al., 2024a) GPT (3.5, 4) Conkas, Oyente, Smartcheck, Honeybadger, Maian, Securify, Slither, Osiris, Smartian, Confuzzius, Sailfish, Mythril, Solhint, AChecker
Ni et al. (Ni et al., 2024) ChatGPT-3.5 Turbo Devign, ReVeal, IVDetect, LineVul, SVulD
@@
In another study, Ni et al. (Ni et al., 2024) compared sequence-based and graph-based models for vulnerability detection. They found that sequence-based models, such as LineVul (Fu and Tantithamthavorn, 2022) and SVulD (Ni et al., 2023), generally outperformed graph-based models like Devign, ReVeal, and IVDetect (Li et al., 2021) in accuracy and stability. However, graph-based models showed strengths in certain types of vulnerabilities like API Abuse and Security Features. The study also noted that LLMs are not as effective for direct use in vulnerability detection, and that their performance is highly dependent on prompt design.
@@
Some studies indicate that combining LLMs with SATs can improve vulnerability detection rates. Li et al. (Li et al., 2024a) observed that LLMs often struggle with complex code reasoning for vulnerability detection and proposed IRIS, a method that combines LLMs with SATs. IRIS uses LLMs to find specific taint specifications related to a vulnerability class (CWE) in third-party library APIs within a project. IRIS improves these specifications with CodeQL to find security vulnerabilities. They evaluated IRIS on a dataset they created with 120 security vulnerabilities found in real-world Java projects, and validated manually. Results showed that IRIS achieved a 35% increase in vulnerability detection compared to using CodeQL alone, while reducing the number of false positives by 80%.
@@
Similarly, Li et al. (Li et al., 2024b) proposed LLift, a framework that combines LLMs with static analysis (using UBITect, an Use Before Initialization (UBI) discovery tool), to improve bug detection in large and complex codebases, with a focus on UBI bugs in the Linux kernel. First, the pipeline uses UBITect to perform a path-insensitive static analysis for scalability, followed by symbolic execution for precision. Then, in the second phase of the pipeline, an LLM is used to address the undecided cases that UBITect cannot handle. The results showed that LLift reduces the complexity of path-sensitive analysis, improving bug detection accuracy without a high rate of false positives. Additionally, LLift was able to identify 4 UBI bugs in the Linux kernel that UBITect could not confirm. In summary, the combination of LLMs and SATs can enhance the detection of vulnerabilities.
5.4. Factors Influencing LLMs Performance